OpenStudy (anonymous):

A circle whose center is at (1, -3) passes through (7, 5). What is the length of the radius of the circle?

OpenStudy (anonymous):

you are being asked to find the distance between the two points do you know how to find that using the distance formula?

OpenStudy (anonymous):

from 1 to 7 is 6 units from -3 to 5 is 8 units the distance is \[d=\sqrt{6^2+8^2}\]

OpenStudy (anonymous):

compute and get \[\sqrt{36+64}=\sqrt{100}=10\]

OpenStudy (anonymous):

thank u